OpenStudy (sugarmochi):

A straight line passes through the origin and has a gradient of -2/3. Find its equation

OpenStudy (imstuck):

First thing you have to realize is that the gradient is the slope of the line. Therefore, the problem is telling you that the slope of the line is -2/3. It passes through the point (0, 0). From there, you can write an equation in slope-intercept form, solving for the "b", which is the y-intercept. Either that or use the point-slope, which I find to be a bit more tedious.

OpenStudy (imstuck):

The general form for a slope-intercept equation is y = mx + b. We have the slope of -2/3 which will go in for "m", and we also have the point (0, 0), which we can use for the x and the y in the general equation. It will be set up like this: 0 = -2/3(0) + b and solve it for b.

OpenStudy (imstuck):

From this we can see that b = 0, so the general form for the equation with a gradient of -2/3 passing through the origin is y = -2/3x

OpenStudy (sugarmochi):

ohh thankyou ! there's another question i need help with- do you mind walking me through another question ?

OpenStudy (imstuck):

Not at all!

OpenStudy (sugarmochi):

 For the following graph, find the equation of the line in the form y = mx + B, and then rewrite the equation in general form.

OpenStudy (imstuck):

Ok, nice and easy. Really! ; )

OpenStudy (imstuck):

Start with the "b", which is the y-intercept. The y-intercept exists where x = 0, or even easier to understand, where the line goes through the y-axis. Our line here goes through the y-axis at -2. So that's the 'b' in our equation.

OpenStudy (imstuck):

So far we have y = mx - 2

OpenStudy (imstuck):

Now we have to find the slope of the line. We can either use the formula or we can use the fact that slope is the rise over the run (the change in y over the change in x) to find the slope. In this case, let's use the formula, so you can see how it works. It will work every time to find a slope.

OpenStudy (imstuck):

\[m=\frac{ y _{2}-y _{1} }{ x _{2} -x _{1}}\]

OpenStudy (imstuck):

Our two coordinates, from the line are (0, -2) and (6, 0), right?

OpenStudy (imstuck):

y_2 is 0 and y_1 is -2, x_2 is 6 and x_1 is 0, so fill them in: \[m=\frac{ 0-(-2) }{ 6-0 }\]

OpenStudy (imstuck):

\[m=\frac{ 0+2 }{ 6-0}=\frac{ 2 }{ 6 }=\frac{ 1 }{ 3 }\]

OpenStudy (imstuck):

So the slope is 1/3. Now we can fill THAT in our equation to get \[y=\frac{ 1 }{ 3 }x -2\]

OpenStudy (imstuck):

And there you have it!

OpenStudy (imstuck):

See how easy?!

OpenStudy (sugarmochi):

thankyou so so so much !

OpenStudy (imstuck):

You bet! Math is my first love!

OpenStudy (skullpatrol):

An equation of a line can be found from a. the slope and y-intercept <-------sub m and b into y = mx + b b. the slope and one point on the line <-------sub m and ( x_0, y_0 ) into y - y_0 = m( x - x_0 ) c. any two points on the line. <-------use the two pionts to calculate m then sub either point into b.